OpenStudy (anonymous):

Can anyone help me with this? The equation below shows the decomposition of lead nitrate. How many grams of oxygen are produced when 23.0 g NO2 is formed? 2Pb(NO3)2(s) → 2PbO(s) + 4NO2(g) + O2 (g

OpenStudy (_.mufasa._):

moles NO2 = 11.5 g / 46.0 g/mol=0.25 moles O2 produced = 0.25/4=0.0625 mass O2 = 0.0625 mol x 32 g/mol=2 g

OpenStudy (anonymous):

woah!...haha can you maybe simplify that? plz? T-T I am so bad at chem

OpenStudy (_.mufasa._):

The answers 2 grams, but teachers usually like when you show your work so just post that.

OpenStudy (anonymous):

Oh! okay. thank you!!!